OpenStudy (anonymous):

Explain how to simplify x(2x + 3x2). You may indicate an exponent in your answer with ^. For example, 3x2 as 3x^2.

OpenStudy (anonymous):

2x^2+3x^3

OpenStudy (anonymous):

how do you explain how you did that?

OpenStudy (anonymous):

the distributive property a(b+c)=ab+ac so x(2x+3x^2)= x*2x+x*3x^2=2x^2+3x^3